### v2

@@ 4.3 Evaluation
Figure 5: Comparison on cross-identity experiments. For a fair comparison, we follow the pre-processing strategy of each method.
Cross-identity experiments.  In Tab. 1 and Tab. 2, we also conduct the cross-identity transfer experiments that transfers the expression and pose of different identity into the source identity. As illustrated in Fig. 5, DPE [42] shows visual artifacts and appearance swap, such as face contours and eye shape, due to the insufficient disentanglement of expression and pose in the motion space. HiDe-NeRF [34] scores the highest identity preservation (CSIM) while un-predictable light changes are involved due to the point-wise deformation field on the tri-plane. Our method can imitate the driving expression without appearance swap and generates a video without video-level artifacts such as light changes and flickers. Please refer to our supplementary videos.
Table 3: Ablation study on the expression encoding. We employ the same evaluation setting with Tab. 1. The best score for each metric is in bold.
@@
Figure 6: Ablation studies on expression encoding. The first row is the results of same-identity transfer, and the second row is the results of cross-identity transfer.
4.4 Ablation Studies
Ablation on the expression encoding.  In Tab. 3, we conduct ablation studies on different expression encoding strategies. In Direct 3DMM, we replace our CLeBS with fully-connected layers to directly inject the expression parameters of 3DMM through EAdaLN. As illustrated in Fig. 6, the direct injection does not change appearance in the same-identity transfer however, it changes appearance (e.g., eyebrows, eyelids, and face contour) as well . Furthermore, since the raw expression parameters inherently contain noise, the generated image exhibits visual artifacts in the cross-identity transfer. In E2E LeBS, we vary the the number of basis vectors $n$ (25, 10, 5) in LeBS to demonstrate the effectiveness of the contrastive pre-training to dicard the appearance information. In these scenarios, LeBS is trained together with entire model in end-to-end manner. Due to the entanglement of appearance and expression, both appearance and expression are changed as a whole as the the number of basis vector $n$ decreases. LeBS alone is insufficient for extracting appearance-free expression from the expression parameters.
Visualization of facial expression parameters.  In Fig. 8, we sample 10 random frames from 10 different videos of distinct people in VFHQ [63] and visualize the low-dimensional t-SNE [39] results of the two expression parameters: $\beta\in\mathbb{R}^{64}$ and $\beta^{\prime}\in\mathbb{R}^{d}$. As depicted in Fig. 7(a), the 3DMM expression parameters exhibit entanglement with their respective source identity, indicating that they still contain appearance information. On the other hand, contrastive pre-training mitigates the entanglement between appearance and expression (Fig. 7(b)), thereby resolving the appearance swap in the cross-identity expression transfer in Fig. 6.
@@ 4.4 Ablation Studies
Figure 8: Linear scaling along the different basis in CLeBS. We visualize the different expression directions along the basis vector in $\mathbf{v}_{4},\mathbf{v}_{8}\in\textbf{V}$.
@@
Novel-view synthesis with expression transfer.  In Fig. 9, we compare the results of novel-view synthesis with expression transfer to those of HiDe-NeRF [34]. Both methods utilize the tri-plane and differentiable volume rendering to generate novel-view images. However, while HiDe-NeRF transfers the driving expression by deforming the generated tri-plane into a canonical tri-plane based on driving conditions [43], our method relies on the hybrid generator $\mathbf{G}$ with EAdaLN. In both the same-identity and cross-identity transfer scenarios, our approach exhibits more view-consistent results, highlighting the effectiveness of our method in expression transfer without relying on deformation. Please refer to supplementary videos.
Figure 9: Novel-view synthesis results with expression transfer.
@@ 5 Conclusion
We presented Export3D, a 3D-aware portrait image animation model that controls the facial expression and the camera view of a source image by leveraging the driving 3DMM expression and camera parameters. Since the expression parameters are still entangled with appearance information, we proposed a contrastive pre-training framework to extract appearance-free expression latent from the parameters. This latent is injected into our generator through expression adaptive layer normalization (EAdaLN) that produces a tri-plane of source identity and driving expression. Finally, differentiable volume rendering renders the tri-plane into 2D images of different views. Extensive experiments show that our contrastive pre-training framework removes the appearance information from the expression parameters, enabling our model to transfer the cross-identity expressions without undesirable appearance swap.

### v3

@@ 4.2 Evaluation
Figure 5: Comparison on cross-identity experiments. For a fair comparison, we follow the pre-processing strategy of each method. Notably, most portrait animation methods fail to preserve the source identity or transfer driving appearance features, such as eye shape and facial contour, in cross-identity scenarios.
@@
Table 3: Ablation studies on the expression encoding. Same evaluation setting with Tab. 1. The best score for each metric is in bold.
@@
Figure 6: Ablation studies on the expression encoding. Without our contrastive pre-training, the expression encoders transfer the expression together with the appearance, such as eyelids and the head size.
Cross-identity experiments.  In Tab. 1 and Tab. 2, we also conduct the cross-identity transfer experiments that transfers the expression and pose of different identity into the source identity. As illustrated in Fig. 5, DPE [42] shows visual artifacts and appearance swap, such as face contours and eye shape, due to the insufficient disentanglement of expression and pose in the motion space. HiDe-NeRF [34] scores the highest identity preservation (CSIM) while un-predictable light changes are involved due to the point-wise deformation field on the tri-plane. Our method can transfer the driving expression without appearance swap and generates a video without video-level artifacts such as light changes and flickers. Please refer to our supplementary videos.
4.3 Ablation Studies and Further Results
Figure 7: Cross-attention vs. EAdaLN.
Table 4: Ablation studies on EAdaLN. The best score for each metric is in bold. We replace EAdaLN in $\mathbf{G}$ with cross-attention to verify the effectiveness of EAdaLN.
Method Same-identity Cross-identity
[…]
Ours (w. Cross-attention) 0.678 0.125 0.042 0.631 0.271 0.122
Ours (w. EAdaLN) 0.811 0.082 0.030 0.694 0.208 0.080
Ablation studies on the expression encoding.  In Tab. 3, we conduct ablation studies on different expression encoding strategies. In Direct 3DMM, we replace our CLeBS with fully-connected layers to directly inject the expression parameters of 3DMM through EAdaLN. As illustrated in Fig. 6, the direct injection does not change appearance when transferring same-identity expression however, it changes appearance (e.g., eyebrows and facial contour) when transferring cross-identity expression. Furthermore, since the raw expression parameters inherently contain noise, the generated image also exhibits visual artifacts. In E2E LeBS, we decrease the the number of basis vectors $n$ in LeBS for appearance bottleneck to validate the proposed contrastive pre-training. Each LeBS with $n=25,10,5$ is jointly trained (i.e., E2E) with entire model without any pre-training. Due to the entanglement of appearance and expression, both appearance and expression are changed as a whole as the the number of basis vector $n$ decreases. LeBS alone is insufficient for extracting appearance-free expression from the expression parameters.
Ablation studies on EAdaLN.  In Tab. 4, we conduct ablation studies on EAdaLN (w. EAdaLN) by comparing it with cross-attention (w. Cross-attention), which is a widely used conditioning method in transformer-based architectures [57, 44]. Specifically, we replace all the EAdaLN blocks in $\mathbf{G}$ (Fig. 3) with cross-attention blocks. In both scenarios, CLeBS serves the refined expression $\beta^{\prime}$. As shown in Tab. 4 and Fig. 7, the cross-attention fails to handle the expression accurately, which verifies the effectiveness of our EAdaLN for the expression conditioning.
Visualization of facial expression parameters.  In Fig. 10, we sample 10 random frames from 10 different videos of distinct individual in VFHQ [64] and visualize the low-dimensional t-SNE [39] results of the two expression parameters: $\beta\in\mathbb{R}^{64}$ and $\beta^{\prime}\in\mathbb{R}^{d}$. In Fig. 8(a), the 3DMM expression parameters show strong entanglement with respective to their identities, indicating the hidden appearance information in them. On the other hand, as shown in Fig. 8(b), our contrastive pre-training mitigates the entanglement, thereby resolving the appearance swap in the cross-identity expression transfer in Fig. 6.
@@ 4.3 Ablation Studies and Further Results
Figure 9: Linear scaling along the different basis vectors of CLeBS. We visualize the different expression directions along the basis vectors $\mathbf{v}_{4},\mathbf{v}_{8}\in\textbf{V}$.
Figure 10: Novel-view synthesis results with expression transfer. Our method can generate more multi-view consistent images compaired to HiDe-NeRF [34].
@@
Novel-view synthesis with expression transfer.  In Fig. 10, we compare the results of novel-view synthesis with expression transfer to those of HiDe-NeRF [34]. Both methods utilize the tri-plane and differentiable volume rendering to generate novel-view images. However, while HiDe-NeRF transfers the driving expression by deforming the generated tri-plane into a canonical tri-plane based on driving conditions [43], our method relies on the hybrid generator $\mathbf{G}$ with EAdaLN. In both same-identity and cross-identity transfer scenarios, our method synthesizes more view-consistent results, highlighting the effectiveness of our method in expression transfer without relying on deformation. Please refer to supplementary videos.
@@ 5 Conclusion
We presented Export3D, a 3D-aware portrait image animation model that controls the facial expression and the camera view of a source image by leveraging the driving 3DMM expression and camera parameters. Since the expression parameters are still entangled with appearance information, we proposed a contrastive pre-training framework to extract appearance-free expressions from the parameters. These refined expressions are injected into our generator through expression adaptive layer normalization (EAdaLN) that produces a tri-plane of source identity and driving expression. Finally, differentiable volume rendering renders the tri-plane into 2D images of different views. Extensive experiments show that our contrastive pre-training framework removes the appearance information from the 3DMM expression parameters, enabling our model to transfer the cross-identity expressions without undesirable appearance swap.